# Decoding Financial Health in Kenya’s Medical Insurance Sector: A Data-Driven Cluster Analysis

### v1

@@ 6   Conclusion
This study identified four distinct clusters of insurance companies based on their financial performance and reporting patterns, utilizing LSTM networks for dimensionality reduction and KMeans clustering. The use of LSTMs enabled the extraction of meaningful temporal patterns, facilitating the grouping of companies with similar trajectories. Clusters 1 and 3 represent companies with relatively stable performance, with Cluster 1 showing resilience despite occasional dips and Cluster 3 reflecting consistent improvements in later stages. In contrast, Clusters 2 and 4 highlight irregular reporting behaviors, where data discontinuities or late reporting pose challenges for assessing long-term financial stability.
While the findings provide valuable insights into the dynamics of the insurance sector, the study also underscores the importance of addressing data gaps and scaling techniques to improve clustering accuracy. Incorporating non-financial metrics and exploring advanced analytical methods could further enhance the depth and breadth of future studies. By addressing these limitations, future research can provide a more nuanced understanding of the factors influencing financial performance and reporting behaviors in the insurance industry.
@@ 11   Funding
12   Acknowledgement
The Stipendium Hungaricum Scholarship Program with Application No. 118250 supported the author.

### v2

@@ 6   Conclusion
This study identified four distinct clusters of insurance companies based on their financial performance and reporting patterns, leveraging LSTM networks for dimensionality reduction and K-Means and Hierarchical clustering for pattern identification. The application of LSTM enabled the extraction of meaningful temporal trends, allowing for the grouping of companies with similar financial trajectories. Clusters 1 and 2 comprise companies with relatively stable performance, with Cluster 2 demonstrating resilience despite occasional fluctuations, while Cluster 1 reflects steady improvements in later stages. In contrast, Clusters 3 and 4 reveal irregular reporting behaviors, where data discontinuities and delayed reporting create challenges in assessing long-term financial stability. While these findings offer valuable insights into the dynamics of the insurance sector, they also highlight the critical need to address data gaps and refine scaling techniques to enhance clustering accuracy. Future research could integrate non-financial metrics and explore advanced analytical methods to deepen the understanding of financial performance and reporting behaviors. By addressing these limitations, future studies can provide a more comprehensive and nuanced perspective on the factors influencing stability and transparency in the insurance industry.